OpenStudy (anonymous):

2radical of 125-2radical of 5 Subtract by collecting like terms.

jimthompson5910 (jim_thompson5910):

\[\Large 2\sqrt{125}-2\sqrt{5}\] \[\Large 2\sqrt{25*5}-2\sqrt{5}\] \[\Large 2\sqrt{25}*\sqrt{5}-2\sqrt{5}\] \[\Large 2*5*\sqrt{5}-2\sqrt{5}\] \[\Large 10\sqrt{5}-2\sqrt{5}\] \[\Large (10-2)\sqrt{5}\] \[\Large 8\sqrt{5}\] So \[\Large 2\sqrt{125}-2\sqrt{5}=8\sqrt{5}\]